OpenStudy (anonymous):

help?? Given f(x) = the quantity of x minus 5, divided by 3, solve for f-1(3).

OpenStudy (anonymous):



OpenStudy (anonymous):

f-1 = 3X+5 f-1(3) = 3(3) +5 = 14